For discharging process

R=10KiloOHM,V=470microF, E=12V.

A)What value will be the voltage across the capacitor 2 time constant

B)what value will be the voltage across after 6 seconds.

C)when the capacitor will be fully charged

D)when will the current through the circuit be reduced by 25%


1
Expert's answer
2021-07-12T12:16:18-0400


A) Time constant

"τ=RC \\

= 10 \;k \Omega\times 470\; \mu F \\

= 4.7 \;s \\

t = 2 \times τ \\

t=9.4 \;s"

Voltage across the capacitor during charging

"V=V_0(1 -e^{-t/RC}) \\

= 12(1 -e^{-9.4/4.7}) \\

= 10.375\;V"

B) Voltage across the capacitor after 6 s

"V=V_0(1 -e^{-t/RC}) \\

= 12(1 -e^{-6/4.7}) \\

= 8.652 \;V"

C) At time t=4τ, voltage V=11.78 V

At time t=5τ, voltage V=11.92 V

At time t=6τ, voltage V=11.97 V

The choice is up to you, which time you want to consider because all time voltage is approximately equal to supply voltage.

D) Current in circuit

"I_c= \frac{V}{R}e^{-t/RC}"

When current reduced by 25%, remaining current 75%:

"0.75 = e^{-t/4.7} \\

ln(0.75)=-t/4.7 \\

t = 1.35 \;s"